OpenStudy (anonymous):

Ben made some home improvements for $9,375 using a six-month deferred payment plan with an interest rate of 19.42%. What is the balance after the deferment period if no payments are made?

OpenStudy (phi):

I think he owes 1/2 a year's interest on top of the original loan

OpenStudy (anonymous):

well the choices are these $10,322.95 $9,375.00 $11,195.63 $10,244.95

OpenStudy (phi):

Do you know how to figure the interest for 1/2 a year at a rate of 19.42% ?

OpenStudy (anonymous):

no im completely lost in this class

OpenStudy (phi):

Did they tell you any formulas ? Does this one ring any bells? http://math.about.com/od/businessmath/ss/Interest.htm

OpenStudy (anonymous):

oh, yes but how would i set this problem up

OpenStudy (phi):

I just checked, and none of the answers match "simple interest"

OpenStudy (phi):

so it looks like a compounded monthly problem

OpenStudy (anonymous):

oh ok thanks

OpenStudy (phi):

Have you seen that formula before?

OpenStudy (phi):

Can you write it down, because people use different symbols for the same formula

OpenStudy (anonymous):

um ok is this how you set it up A=9375(1+19.42%/12)^12

OpenStudy (phi):

before you put in numbers,what are the letters you use?

OpenStudy (anonymous):

a=p(1+r/n)nt

OpenStudy (phi):

OK except you mean a=p(1+r/n)^(nt) the ^ means raise to the power (otherwise you might think multiply, which is not right) You are on the right track. First, how many months is this loan ? Read the problem carefully.

OpenStudy (anonymous):

oh six months

OpenStudy (phi):

that means t = 1/2 of a year n is the number of times compounded in a year. n is 12 the number nt should be 6: 12*1/2 = 6 generally I just think : how many months is the loan so we know nt = 6 next, the rate is ALWAYS CHANGED INTO A DECIMAL % means divide by 100

OpenStudy (phi):

19.42% means 19.42 per cent (and cent is latin for hundred. think dollar and cents or century) 19.42% = 19.42/100

OpenStudy (anonymous):

0.1942

OpenStudy (phi):

ok, now what do you have when you put in the numbers?

OpenStudy (anonymous):

A=9375(1+0.1942%/6)^6

OpenStudy (phi):

almost, by n is 12. This loan is going to be paid monthly, so you divide the interest rate (which is for a whole year) by 12. But the loan is for 6 months. also, once you change 19.42% to a decimal it is not a percent anymore. (no % any more) can you fix your formula?

OpenStudy (anonymous):

im lost.....is not A=9375(1+0.1942%/12)^6 if not can you set it up?

OpenStudy (phi):

that looks good except for the % A=9375(1+0.1942/12)^6

OpenStudy (anonymous):

oh ok sorry =D

OpenStudy (anonymous):

10322?

OpenStudy (phi):

now copy and paste into google 9375(1+0.1942/12)^6= or get your calculator

OpenStudy (phi):

I got 10322.9466 round to the nearest penny. That means look at the decimal to the right of 94 and check if it is 5 or more. if it is, round up to 95 cents (otherwise, just make it 94 cents) in this case, we get 10322.95